A little girl throws a Frisbee. It is 3 feet above the ground at the time she lets go. The Frisbee follows a parabolic path and after 5 seconds it is 6 feet above the ground. After 10 seconds it is again six feet above the ground.
a. Write the particular equation of the quadratic
b. How high is the frisbee after 3 seconds?
c. How high is the frisbee after 12 seconds?

Answer by stanbon(75887)

A little girl throws a Frisbee. It is 3 feet above the ground at the time she lets go. The Frisbee follows a parabolic path and after 5 seconds it is 6 feet above the ground. After 10 seconds it is again six feet above the ground.
----
You have three points: (0,3),(5,6),(10,6)
---------------
a. Write the particular equation of the quadratic
Substitute each point-pair into the following form
to get three equations with variables a,b,c.
============
Form ax^2 + bx + c = y
0 + 0 + c = 3
25a + 5b + c = 6
100a+10b + c = 6
-------------------
Solve by any method you know to get:
f(x) = -0.06x^2 + 0.9x + 3
---------------------------
b. How high is the frisbee after 3 seconds?
f(3) = 5.16 ft.
------
c. How high is the frisbee after 12 seconds?
f(12) = 5.16 ft.
=======================
Cheers,
stan H.